U.S. Robotics Inc. has a current capital structure of 30% debt and 70% equity. Its current before-tax cost of debt is 8%, and its tax rate is 35%. It currently has a levered beta of 1.15. The risk-free rate is 3.5%, and the risk premium on the market is 7.5%. U.S. Robotics Inc. is considering changing its capital structure to 60% debt and 40% equity. Increasing the firms level of debt will cause its before-tax cost of debt to increase to 10%. First, solve for U.s. Robotics Inc.s unlevered beta. Relever U.S. Robotics Inc.s beta using the firms new capital structure. Use U.S. Robotics Inc.s levered beta under the new capital structure, to solve for its cost of equity under the new capital structure. What will the firms weighted average cost of capital (WACC) be if it makes this change in its capital structure? o 7.5% o 10.7% O 8.090 10.2%
Answer #1

unlevered Beta=1.15/(1+(1-35%)*(30%/70%))=0.90

2) new Beta=0.90*(1+(1-35%)*(60%/40%))=1.78

3)

cost of equity=3.5%+1.78*7.5%=16.9%

4)

WACC=40%*16.9%+60%*10%*(1-35%)=10.7%

the above is answer..